# SAT: Dynamic Spatial Aptitude Training for Multimodal Language Models

### v2

@@ Abstract
Reasoning about motion and space is a fundamental cognitive capability that is required by multiple real-world applications. While many studies highlight that large multimodal language models (MLMs) struggle to reason about space, they only focus on static spatial relationships, and not dynamic awareness of motion and space—i.e., reasoning about the effect of egocentric and object motions on spatial relationships. Manually annotating such object and camera movements is expensive. Hence, we introduce SAT, a simulated spatial aptitude training dataset comprising both static and dynamic spatial reasoning across 175K question-answer (QA) pairs and 20K scenes. Complementing this, we also construct a small (150 image-QAs) yet challenging dynamic spatial test set using real-world images. Leveraging our SAT datasets and 6 existing static spatial benchmarks, we systematically investigate what improves both static and dynamic spatial awareness. Our results reveal that simulations are surprisingly effective at imparting spatial aptitude to MLMs that translate to real images. We show that perfect annotations in simulation are more effective than existing approaches of pseudo-annotating real images. For instance, SAT training improves a LLaVA-13B model by an average $11\%$ and a LLaVA-Video-7B model by an average $8\%$ on multiple spatial benchmarks, including our real-image dynamic test set and spatial reasoning on long videos—even outperforming some large proprietary models. While reasoning over static relationships improves with synthetic training data, there is still considerable room for improvement for dynamic reasoning questions.
@@ 1 Introduction
Reasoning about space and motion is a fundamental cognitive capability that allows humans and animals to survive and operate in the real world (Gardner, 2011; Vasilyeva & Lourenco, 2012; Blades & Spencer, 1994). While simple visual relationships (such as left-right) may be trivial to recognize for humans, more dynamic reasoning involving motion such as taking perspectives and reasoning about the effect of ego-motion or object movement requires more cognitive support (Gardner, 2011; Tversky, 2019). Hence, it is not surprising that such “dynamic” spatial reasoning capabilities in children correlate with their aptitude in geometry, physics, and even linguistic reasoning (Taylor & Tversky, 1992; Mallot, 2024).
@@
Despite the promise of multimodal language models (MLMs) as general-purpose intelligent agents (Achiam et al., 2023; Zhu et al., 2023; Liu et al., 2024), recent studies reveal that these models still struggle with spatial reasoning (Chen et al., 2024; Cheng et al., 2024; Tong et al., 2024; Fu et al., 2024b). Specifically, MLMs often fail to predict the relative positions of objects in static images—a limitation attributed to the scarcity of spatial relationship annotations in their training data. To address this, recent methods (Cheng et al., 2024; Chen et al., 2024) introduce pseudo-annotations to encode static object relationships in real images (Cheng et al., 2024), which require extensive engineering and refining to scale. Furthermore, many real-world applications require reasoning that extends beyond static object positions. For example, smart glasses and embodied AI applications need to reason about dynamic scenes, where reasoning about the effect of object and camera movements are essential (Duan et al., 2024; Yuan et al., 2024a; Liu et al., 2025). Generating pseudo-annotations using existing methods for such dynamic settings is non-trivial due to the complexity of predicting action causality on real images.
@@
We explore the possibility of using simulation data as a simple solution to impart both static and dynamic spatial reasoning into MLMs. We propose Spatial Aptitude Training (SAT), an approach to generate spatial question-answer (QA) data without any human supervision. Manually annotating 3D movements is expensive; hence, SAT leverages 22K ProcTHOR (Deitke et al., 2022) scenes composed of 1K assets to generate 175K QA pairs. With perfect 3D information and control of the assets, SAT goes beyond static object relationships to questions that require reasoning about egocentric and object movements. Since our data is generated procedurally by composing assets, it can be scaled up arbitrarily. Hence, SAT is programmatically controllable, unlike existing 3D (Brazil et al., 2023; Azuma et al., 2022) or spatial datasets (Cheng et al., 2024), which are also not compositional.
@@
With SAT, we analyze what types of training data improve spatial reasoning. We focus on two types of spatial reasoning data. First, spatial-QA about object relations in static scenes, shown in Fig. 1 (under “Existing Spatial Bench”), to impart reasoning about the relative locations of objects in the scene (e.g., where is object X with respect to object Y?). Next, we evaluate the effect of dynamic spatial tasks, Fig. 1 (shown under “Our SAT Dynamic Bench”). This includes questions about egocentric movement, object movement, allocentric perspective, goal aiming, and action consequences. These dynamic spatial reasoning skills are well-studied in human cognitive development: children understand the consequences of self-motion (e.g., the moving room test (Anderson et al., 2013)), track how entities move in a scene (Anderson et al., 2013), and can adopt others’ viewpoints when planning actions (Brucato et al., 2023; Vasilyeva & Lourenco, 2012; Blades & Spencer, 1994).
@@
We use two widely adopted open-source MLMs, an image-based LLaVA-1.5-13B and a video-based LLaVA-Video-7B, as our base models for evaluations. To test static spatial reasoning, we use four contemporary real image benchmarks: CV-Bench (Tong et al., 2024), BLINK (Fu et al., 2024b), Visual Spatial Relations (VSR) dataset (Liu et al., 2023), and GQA-Spatial (Hudson & Manning, 2019a). In addition, we test the effect of SAT on spatial reasoning on longer videos using the recent VSI-Bench (Yang et al., 2024a). Since no dataset exists for dynamic spatial reasoning, we construct a difficult SAT test set on real images.
@@
Challenging conventional wisdom about sim-to-real transfer in model training, we find simulations to be surprisingly effective at imparting spatial intelligence to MLMs. Our simulated SAT training improves the baseline LLaVA-13B model by $11\%$ and LLaVA-video-7B by $8\%$ (absolute) on average on a wide range of spatial benchmarks, including our challenging dynamic test set—even outperforming some larger closed-source and spatially-tuned models (Achiam et al., 2023; Team et al., 2023). Interestingly, SAT also improves spatial reasoning on long videos (Yang et al., 2024a)—especially in route planning by $4\%$, showing promise for aiding embodied applications. However, while simple static relationships in existing datasets are easy to improve, our best model performs at $62\%$ on our SAT dynamic questions (random chance is $50\%$)—hence, leaving considerable room for improvement.
@@ 2 Related work
Our work draws inspiration from fundamental schools of thought in neuroscience that suggest spatial intelligence is a core foundation for most cognitive abilities (Tversky & Suwa, 2009; Mallot, 2024; Duan et al., 2022b).
@@
Spatial understanding benchmarks. We visualize a comparison of our dataset with existing 3D and spatial benchmarks (Table 1). 3D understanding is vital for various computer vision tasks, including segmentation (Jatavallabhula et al., 2023; Hong et al., 2023a; Kerbl et al., 2023; Nie et al., 2020; Kundu et al., 2022; Wang, 2023), object localization (Qi et al., 2019; Sajjadi et al., 2022; Kamath et al., 2021), tracking (Agia et al., 2022; Kurenkov et al., 2023; Li et al., 2022), fine-grained scene captioning (Chen et al., 2021b; Su et al., 2021), open-vocabulary classification (Chen et al., 2020; Hong et al., 2023a; Achlioptas et al., 2020; Shao et al., 2019), and question answering (Ye et al., 2021; Azuma et al., 2022; Linghu et al., 2024). Most use 3D scans of real-world data (Hong et al., 2023b), but 3d scans are expensive to recompute for dynamic scenes and most MLMs are trained to input 2D images-hence we focus on 3D spatial reasoning on 2D inputs. In light of potential applications in embodied AI (Duan et al., 2022b), we align our dataset to be compatible with physics simulators, allowing us to go beyond datasets which only leverage image annotations (Caesar et al., 2020) and have static spatial reasoning (Fu et al., 2024b; Duan et al., 2022a; Chen et al., 2024; Kamath et al., 2023; Su et al., 2021; Liu et al., 2023; Goyal et al., 2020; Yang et al., 2019; Du et al., 2024).
@@
Synthetic training data. Many have investigated whether synthetic information can boost reasoning in real environments Geng et al. (2024), finding that synthetic images sometimes help in classification (Chen et al., 2021a), semantic understanding (Mishra et al., 2022), correcting biases (Qraitem et al., 2023), and teaching navigation capabilities to embodied AI (Ehsani et al., 2024; Silwal et al., 2024). Inspired by this, our work explores if synthetic data with perfect 2D/3D information can improve spatial reasoning.
@@
Vision-language models. Our task is heavily influenced by the emergence of multimodal foundation models (Radford et al., 2021; Jia et al., 2021; Wang et al., 2022a; Yuan et al., 2021; Fu et al., 2021; Wang et al., 2022b; Ye et al., 2023). VLMs have been adapted for a wide range of downstream image (Tsimpoukelli et al., 2021; Driess et al., 2023; Zhu et al., 2023; Li et al., 2022; Alayrac et al., 2022; Dai et al., 2024) and video (Zhang et al., 2023; Maaz et al., 2023; Tan et al., 2024; Wang et al., 2024; Li et al., 2023; Luo et al., 2023) understanding tasks. Many tasks require spatial understanding of the scene (Brohan et al., 2023; Zhang et al., 2024b). Multiple recent works (Fu et al., 2024a; Hong et al., 2023a; Cho et al., 2024) have noted their weakness in 3D pose and location estimation. Adjacent to works like Ray et al. (2023); Zhang et al. (2024a); Hsieh et al. (2023); Thrush et al. (2022) that point to deficiencies in understanding compositions of spatial relationships with objects and attributes, we explore if perfect 3D information in synthetic images can improve spatial understanding.
@@ 3 SAT: Spatial Aptitude Training
In total, our training dataset contains $175$K questions across $22$K procedurally generated scenes from ProcTHOR-10K dataset of indoor apartment buildings. Next, we outline the generation process for the two kinds of spatial questions—static and dynamic.
@@
Aligning with contemporary benchmarks for spatial reasoning, we first generate instruction-tuning data for static spatial reasoning that deal with relative relations of objects. Overall, we generate $127$K static spatial QA pairs across $8$K images across the following types:
@@
Count. Since many MLMs (Goyal et al., 2017) struggle with counting, we include counting questions using the number of object instances from the metadata of our simulator.
@@ 3.2 Generating Dynamic Spatial QA
Grounded in spatial cognitive tests (Anderson et al., 2013; Brucato et al., 2023; Vasilyeva & Lourenco, 2012), we outline five different complex tasks that require reasoning about egocentric movements, object movements, and allocentric perspectives. We generate $48$K QAs on $13$K images. The high-level idea behind generating such QAs is illustrated in Figure 2. Given a frame in a simulated environment, we take an action and formulate QAs based on how the 3D orientation of objects changes based on the action taken. Below, we outline the specific approach for each of the question types.
@@
Egocentric Movement. This is based on the “moving room test” Anderson et al. (2013), a fundamental test designed to assess and improve spatial cognitive development in children. This test aims to measure if an agent can judge how they moved given two frames. This is useful beyond just measuring spatial cognition since this task can help judge navigation trajectories from egocentric videos in embodied applications. We take a random action from the choices of rotating left or right by a random angle, or moving left/right by a random distance. Note that moving and rotating are different since a camera can be moving left while rotating right. Based on the action taken, we formulate a question of the type: How did the camera taking the video likely rotate/move? with the answer being the action sequence taken. We have $6.9$K training image-QA pairs of this type. This is denoted as EgoM in the tables. More details of the exact simulator actions taken to generate this are in the appendix.
@@
Object Movement. Here, we randomly choose an object and move it by a randomly chosen distance and direction, ensuring that the object is still in the frame of view. Next, we compare the updated 3D position of the object with the original position normalized by the camera coordinates to decide if the object got closer, further, more left, or more right from the original position, or if it did not move. Based on that, we form QA pairs of the type: Did any of the objects move from the first frame to the second frame? with the answer being the way the object moved. Note that sometimes objects may move in conjunction with camera movement. To answer the questions accurately, the agent needs to learn to distinguish between egocentric movement and objects moving. We have $6.9$K training image-QA pairs of this type. This is denoted as ObjM in the tables.
@@
Allocentric Perspective. Inspired by a spatial cognitive test for humans and animals Brucato et al. (2023), this test checks if the agent is able to take the perspective of another viewer and judge the relative locations of objects according to the other viewer. To make such reasoning QAs, we first choose a 2D point in the scene and mark it as “X". Next, we teleport that agent to the 3D location corresponding to “X" (determined by ray tracing). We check the relative positions of objects according to the camera view from “X” (similarly as described in 3.1. We make QA’s of the type: For someone at the mark ‘X’ facing left/right by 90 degrees, would the <object> be to their left or right? We have $6$K training image-QA pairs of this type. This is denoted as Pers in the tables.
@@
Goal Aiming. Aiming is a prerequisite for efficient navigation to objects Franz & Mallot (2000), a fundamental spatial cognitive capability. Hence, we design QAs that check how well agents can aim to the desired object. We pick a random object and calculate the angle of the object to the camera using the 3D location of the object and camera, assuming looking forward is 0 degrees (exact equations in the supplementary). Based on the angle, we formulate questions of the type: I need to go to the countertop. Which direction should I turn to face it? Since precise angles are hard to judge from a single image, we formulate answers as choices of rough angles to turn towards the left or right. We have $6.8$K training image-QA pairs of this type. This is denoted as Aim in the tables.
@@
Action Consequence. Here, the agent needs to reason about how the spatial relationships change when it takes a movement action, inspired by how humans can reason about the consequences of their movements in an environment Franz & Mallot (2000). Here, we show the first frame and ask the agent to judge if we would move closer/further, or look towards or away from an object if it took that action. e.g., If I rotate left by 90 degrees and move forward, would I move further from the sofa? Note that in most cases, moving forward would get us closer to an object. To make the distribution of answers even, we sometimes rephrase the question as to whether we would be facing the object or not. We have $15$K training image-QA pairs of this type. This is denoted as EgoAct in the tables.
@@
Real-image Test Set. Existing image spatial datasets only test for static relationships and not the above-mentioned dynamic capabilities. A recent video spatial dataset (Yang et al., 2024a) also focuses on static object relations in a video input where the scene doesn’t change. Hence, we construct a small yet challenging test set of dynamic QAs that focus on how relationships change given ego or object motions and varying perspectives. We use expert annotators to annotate 150 image-question-answer pairs ($\sim 30$ each of the above dynamic types). We also generate a larger synthetic test set—$647$ object movement, $647$ egocentric movement, $592$ goal aim, $1336$ action consequence, and $778$ perspective questions on $805$ images for analysis on individual splits.

### v3

@@ Abstract
Reasoning about motion and space is a fundamental cognitive capability that is required by multiple real-world applications. While many studies highlight that large multimodal language models (MLMs) struggle to reason about space, they only focus on static spatial relationships, and not dynamic awareness of motion and space—i.e., reasoning about the effect of egocentric and object motions on spatial relationships. Manually annotating such object and camera movements is expensive. Hence, we introduce SAT, a simulated spatial aptitude training dataset utilizing 3D simulators, comprising both static and dynamic spatial reasoning across 175,000 question-answer (QA) pairs and 20,000 scenes. Complementing this, we also construct a small (150 image-QAs) yet challenging dynamic spatial test set using real-world images. Leveraging our SAT datasets and 6 existing static spatial benchmarks, we systematically investigate what improves both static and dynamic spatial awareness. Our results reveal that simulations are surprisingly effective at imparting spatial aptitude to MLMs that translate to real images. We show that perfect annotations in simulation are more effective than existing approaches of pseudo-annotating real images. For instance, SAT training improves a LLaVA-13B model by an average $11\%$ and a LLaVA-Video-7B model by an average $8\%$ on multiple spatial benchmarks, including our real-image dynamic test set and spatial reasoning on long videos—even outperforming some large proprietary models. While reasoning over static relationships improves with synthetic training data, there is still considerable room for improvement for dynamic reasoning questions.
@@ 1 Introduction
Reasoning about space and motion is a fundamental cognitive capability that allows humans and animals to survive and operate in the real world (gardner2011frames; vasilyeva2012development; blades1994development). While simple visual relationships (such as left-right) may be trivial to recognize for humans, more dynamic reasoning involving motion such as taking perspectives and reasoning about the effect of ego-motion or object movement requires more cognitive support (gardner2011frames; tversky2019mind). Hence, it is not surprising that such “dynamic” spatial reasoning capabilities in children correlate with their aptitude in geometry, physics, and even linguistic reasoning (taylor1992spatial; mallot2024geometry).
@@
Despite the promise of multimodal language models (MLMs) as general-purpose intelligent agents (achiam2023gpt; zhu2023minigpt; liu2024visual), recent studies reveal that these models still struggle with spatial reasoning (chen2024spatialvlm; cheng2024spatialrgptgroundedspatialreasoning; tong2024cambrian1fullyopenvisioncentric; fu2024blink). Specifically, MLMs often fail to predict the relative positions of objects in static images—a limitation attributed to the scarcity of spatial relationship annotations in their training data. To address this, recent methods (cheng2024spatialrgptgroundedspatialreasoning; chen2024spatialvlm) introduce pseudo-annotations to encode static object relationships in real images (cheng2024spatialrgptgroundedspatialreasoning), which require extensive engineering and refining to scale. Furthermore, many real-world applications require reasoning that extends beyond static object positions. For example, smart glasses and embodied AI applications need to reason about dynamic scenes, where reasoning about the effect of object and camera movements are essential (duan2024manipulate; yuan2024robopoint; liu2024coarse). Generating pseudo-annotations using existing methods for such dynamic settings is non-trivial due to the complexity of predicting action causality on real images.
@@
We explore the possibility of using simulation data as a simple solution to impart both static and dynamic spatial reasoning into MLMs. We propose Spatial Aptitude Training (SAT), an approach to generate spatial question-answer (QA) data without any human supervision. Manually annotating 3D movements is expensive; hence, SAT leverages 22K ProcTHOR (deitke2022️) scenes composed of 1K assets to generate 175K QA pairs. With perfect 3D information and control of the assets, SAT goes beyond static object relationships to questions that require reasoning about egocentric and object movements. Since our data is generated procedurally by composing assets, it can be scaled up arbitrarily. Hence, SAT is programmatically controllable, unlike existing 3D (brazil2023omni3d; azuma2022scanqa) or spatial datasets (cheng2024spatialrgptgroundedspatialreasoning), which are also not compositional.
@@
With SAT, we analyze what types of training data improve spatial reasoning. We focus on two types of spatial reasoning data. First, spatial-QA about object relations in static scenes, shown in Fig. 1 (under “Existing Spatial Bench”), to impart reasoning about the relative locations of objects in the scene (e.g., where is object X with respect to object Y?). Next, we evaluate the effect of dynamic spatial tasks, Fig. 1 (shown under “Our SAT Dynamic Bench”). This includes questions about egocentric movement, object movement, allocentric perspective, goal aiming, and action consequences. These dynamic spatial reasoning skills are well-studied in human cognitive development: children understand the consequences of self-motion (e.g., the moving room test (anderson2013role)), track how entities move in a scene (anderson2013role), and can adopt others’ viewpoints when planning actions (brucato2023measuring; vasilyeva2012development; blades1994development).
@@
We use two widely adopted open-source MLMs, an image-based LLaVA-1.5-13B and a video-based LLaVA-Video-7B, as our base models for evaluations. To test static spatial reasoning, we use four contemporary real image benchmarks: CV-Bench (tong2024cambrian1fullyopenvisioncentric), BLINK (fu2024blink), Visual Spatial Relations (VSR) dataset (liu2023visualspatialreasoning), and GQA-Spatial (hudson2019gqanewdatasetrealworld). In addition, we test the effect of SAT on spatial reasoning on videos using the recent VSI-Bench (yang2024thinking) benchmark and MME-RealWorld (zhang2024mme), which include out-of-domain outdoor videos in realistic self-driving scenarios. Since no dataset exists for dynamic spatial reasoning on images (predicting action causality on images), we construct a difficult SAT test set on real images.
@@
Challenging conventional wisdom about sim-to-real transfer in model training, we find simulations to be surprisingly effective at imparting spatial intelligence to MLMs. Our simulated SAT training improves the baseline LLaVA-13B model by $11\%$ and LLaVA-video-7B by $8\%$ (absolute) on average on a wide range of spatial benchmarks, including our challenging dynamic test set—even outperforming some larger closed-source and spatially-tuned models (achiam2023gpt; team2023gemini). Interestingly, SAT also improves spatial reasoning on long videos (yang2024thinking)—especially in route planning by $4\%$, showing promise for aiding embodied applications. However, while simple static relationships in existing datasets are easy to improve, our best model performs at $62\%$ on our SAT dynamic questions (random chance is $50\%$)—hence, leaving considerable room for improvement.
@@ 2 Related work
Our work draws inspiration from fundamental schools of thought in neuroscience that suggest spatial intelligence is a core foundation for most cognitive abilities (tversky2009thinking; mallot2024geometry; duan2022survey).
@@
Spatial understanding benchmarks. We visualize a comparison of our dataset with existing 3D and spatial benchmarks (Table 1). 3D understanding is vital for various computer vision tasks, including segmentation (jatavallabhula2023conceptfusion; hong20233d; kerbl20233d; nie2020total3dunderstanding; kundu2022panoptic; wang2023octformer), object localization (qi2019deep; sajjadi2022object; kamath2021mdetrmodulateddetection), tracking (agia2022taskography; kurenkov2023modeling; li2022pre), fine-grained scene captioning (chen2021scan2cap; su202125dvisualrelationshipdetection), open-vocabulary classification (chen2020scanrefer; hong20233d; achlioptas2020referit3d; shao2019objects365), and question answering (ye20213d; azuma2022scanqa; linghu2024multimodalsituatedreasoning3d). Most use 3D scans of real-world data (hong20233dllm), but 3d scans are expensive to recompute for dynamic scenes and most MLMs are trained to input 2D images-hence we focus on 3D spatial reasoning on 2D inputs. In light of potential applications in embodied AI (duan2022survey), we align our dataset to be compatible with physics simulators, allowing us to go beyond datasets which only leverage image annotations (caesar2020nuscenesmultimodaldatasetautonomous) and have static spatial reasoning (fu2024blink; duan2022pip; chen2024spatialvlm; kamath2023whatsupvisionlanguagemodels; su202125dvisualrelationshipdetection; liu2023visualspatialreasoning; goyal2020rel3d; yang2019spatialsenseadversariallycrowdsourcedbenchmark; du2024embspatialbenchbenchmarkingspatialunderstanding).
@@
Synthetic training data. Many have investigated whether synthetic information can boost reasoning in real environments geng2024unmet, finding that synthetic images sometimes help in classification (chen2021contrastivesyntorealgeneralization), semantic understanding (mishra2022task2sim), correcting biases (qraitem2023fake), and teaching navigation capabilities to embodied AI (ehsani2024spoc; silwal2024we). Inspired by this, our work explores if synthetic data with perfect 2D/3D information can improve spatial reasoning.
@@
Vision-language models. Our task is heavily influenced by the emergence of multimodal foundation models (radford2021learning; jia2021scaling; wang2022omnivl; yuan2021florence; fu2021violet; wang2022internvideo; ye2023mplug). VLMs have been adapted for a wide range of downstream image (tsimpoukelli2021multimodal; driess2023palm; zhu2023minigpt; li2022pre; alayrac2022flamingo; dai2024instructblip) and video (zhang2023video; maaz2023video; tan2024koala; wang2024internvideo2; li2023videochat; luo2023valley) understanding tasks. Many tasks require spatial understanding of the scene (brohan2023rt; zhang2024feedback). Multiple recent works (fu2024scene; hong20233d; cho2024language) have noted their weakness in 3D pose and location estimation. Adjacent to works like ray2023cola; zhang2024task; hsieh2023sugarcrepe; thrush_winoground_2022 that point to deficiencies in understanding compositions of spatial relationships with objects and attributes, we explore if perfect 3D information in synthetic images can improve spatial understanding.
@@ 3 SAT: Spatial Aptitude Training
In total, our training dataset contains 175175K questions across 2222K procedurally generated scenes from ProcTHOR-10K dataset of indoor apartment buildings. Next, we outline the generation process for the two kinds of spatial questions—static and dynamic.
@@
Aligning with contemporary benchmarks for spatial reasoning, we first generate instruction-tuning data for static spatial reasoning that deal with relative relations of objects. Overall, we generate 127127K static spatial QA pairs across 88K images across the following types:
@@
Count. Since many MLMs (goyal2017makingvvqamatter) struggle with counting, we include counting questions using the number of object instances from the metadata of our simulator.
@@ 3.2 Generating Dynamic Spatial QA
Grounded in spatial cognitive tests (anderson2013role; brucato2023measuring; vasilyeva2012development), we outline five different complex tasks that require reasoning about egocentric movements, object movements, and allocentric perspectives. We generate 4848K QAs on 1313K images. The high-level idea behind generating such QAs is illustrated in Figure 2. Given a frame in a simulated environment, we take an action and formulate QAs based on how the 3D orientation of objects changes based on the action taken. Below, we outline the specific approach for each of the question types.
@@
Egocentric Movement. This is based on the “moving room test” anderson2013role, a fundamental test designed to assess and improve spatial cognitive development in children. This test aims to measure if an agent can judge how they moved given two frames. This is useful beyond just measuring spatial cognition since this task can help judge navigation trajectories from egocentric videos in embodied applications. We take a random action from the choices of rotating left or right by a random angle, or moving left/right by a random distance. Note that moving and rotating are different since a camera can be moving left while rotating right. Based on the action taken, we formulate a question of the type: How did the camera taking the video likely rotate/move? with the answer being the action sequence taken. We have 6.96.9K training image-QA pairs of this type. This is denoted as EgoM in the tables. More details of the exact simulator actions taken to generate this are in the appendix.
@@
Object Movement. Here, we randomly choose an object and move it by a randomly chosen distance and direction, ensuring that the object is still in the frame of view. Next, we compare the updated 3D position of the object with the original position normalized by the camera coordinates to decide if the object got closer, further, more left, or more right from the original position, or if it did not move. Based on that, we form QA pairs of the type: Did any of the objects move from the first frame to the second frame? with the answer being the way the object moved. Note that sometimes objects may move in conjunction with camera movement. To answer the questions accurately, the agent needs to learn to distinguish between egocentric movement and objects moving. We have 6.96.9K training image-QA pairs of this type. This is denoted as ObjM in the tables.
@@
Allocentric Perspective. Inspired by a spatial cognitive test for humans and animals brucato2023measuring, this test checks if the agent is able to take the perspective of another viewer and judge the relative locations of objects according to the other viewer. To make such reasoning QAs, we first choose a 2D point in the scene and mark it as “X". Next, we teleport that agent to the 3D location corresponding to “X" (determined by ray tracing). We check the relative positions of objects according to the camera view from “X” (similarly as described in 3.1. We make QA’s of the type: For someone at the mark ‘X’ facing left/right by 90 degrees, would the <object> be to their left or right? We have 66K training image-QA pairs of this type. This is denoted as Pers in the tables.
@@
Goal Aiming. Aiming is a prerequisite for efficient navigation to objects franz2000biomimetic, a fundamental spatial cognitive capability. Hence, we design QAs that check how well agents can aim to the desired object. We pick a random object and calculate the angle of the object to the camera using the 3D location of the object and camera, assuming looking forward is 0 degrees (exact equations in the supplementary). Based on the angle, we formulate questions of the type: I need to go to the countertop. Which direction should I turn to face it? Since precise angles are hard to judge from a single image, we formulate answers as choices of rough angles to turn towards the left or right. We have 6.86.8K training image-QA pairs of this type. This is denoted as Aim in the tables.
@@
Action Consequence. Here, the agent needs to reason about how the spatial relationships change when it takes a movement action, inspired by how humans can reason about the consequences of their movements in an environment franz2000biomimetic. Here, we show the first frame and ask the agent to judge if we would move closer/further, or look towards or away from an object if it took that action. e.g., If I rotate left by 90 degrees and move forward, would I move further from the sofa? Note that in most cases, moving forward would get us closer to an object. To make the distribution of answers even, we sometimes rephrase the question as to whether we would be facing the object or not. We have 1515K training image-QA pairs of this type. This is denoted as EgoAct in the tables.
@@
Real-image Test Set. Existing image spatial datasets only test for static relationships and not the above-mentioned dynamic capabilities. A recent video spatial dataset (yang2024thinking) also focuses on static object relations in a video input where the scene doesn’t change. Hence, we construct a small yet challenging test set of dynamic QAs that focus on how relationships change given ego or object motions and varying perspectives. We use expert annotators to annotate 150 image-question-answer pairs ($\sim 30$ each of the above dynamic types). We also generate a larger synthetic test set—$647$ object movement, $647$ egocentric movement, $592$ goal aim, $1336$ action consequence, and $778$ perspective questions on $805$ images for analysis on individual splits.